The perimeter of a rectangle is 110 inches. The length exceeds the width by 37 inches. Find the length and the width.
Answer by simranb(63)

Let the width be x
then,
Perimeter=2(l+b)=110 -eq 1
perimeter=110
width=x
length=x+37
substitute these values to eq 1
2(x+x+37)=110
2(2x+37)=110
4x+74=110
4x=110-74
4x=46
x=46/4
x=11.5 inches
so,width=11.5 inches
Length=11.5+37=48.5 inches
Hope this helps!